By installing a $120 thermostat that reduces the temperature setting at night, a family hopes to cut its annual bill for heating oil by 8 percent, and thereby recover the cost of the thermostat in fuel savings after 2 years. What was the family's annual fuel bill before installing the thermostat?
Answer by mananth(16946)

By installing a $120 thermostat that reduces the temperature setting at night, a family hopes to cut its annual bill for heating oil by 8 percent, and thereby recover the cost of the thermostat in fuel savings after 2 years. What was the family's annual fuel bill before installing the thermostat?

Let the annual bill for oil be x
reduction is 8/100 *x
for 2 years the reduction will be 8x/100 * 2
8x/100 * 2 = 120
16x/100 =120
16x=12000
x=12000/16
= 750
annual oil bill was $750